# Dissecting Open Edge Computing Platforms: Ecosystem, Usage, and Security Risks

### v2

@@ I Introduction
However, these OECPs are black boxes since little is publicly available regarding their technical mechanisms or real-world usage. Furthermore, it is intuitive to wonder what practical security/privacy implications these OECPs may render, especially considering the claimed extensive adoption by online services. For instance, our preliminary investigation reveals that none of these OECPs enforce strict vetting for edge node operators, and any third party, can turn their devices into edge nodes, which can introduce byzantine or even malicious edge nodes. It also breaks the traditional threat model for edge computing wherein edge nodes are fully controlled by the ECP and are usually considered as trusted. To clear away the dense mist surrounding OECPs, we conduct an extensive security study on two representative ones, namely, TipTime [10] and OneThingCloud [9]. As elaborated in §II, these two stand out in a comparative analysis of all identified OECPs. Our study would not be possible without the design and implementation of a set of automatic or semi-automatic tools. Particularly, to capture edge activities and edge traffic, a deployment framework is designed to deploy edge nodes across OECPs at different locations and capture their edge traffic. Then, given edge traffic of multiple TBs, an efficient edge traffic analyzer is built up to concurrently process edge traffic flows, group them into pre-defined traffic categories, as well as correlating them with various real-world parties. Also, a security testbed is carefully designed to empirically evaluate security risks of OECPs while avoiding interfering with real-world edge computing activities. Applying these tools leads to new understandings of OECPs with regards to their technical mechanisms, the landscape of edge nodes, real-world usage, as well as security and privacy risks. Our key findings are highlighted as below.
@@
First of all, Both OECPs share many properties in their technical mechanism and service model. Briefly, both enable the co-deployment of different edge computing tasks on a single edge node and each edge computing task runs like an independent kingdom inside the edge node. Also, when it comes to edge node recruitment, both support a variety of computing devices including mobile phones, desktop/laptop computers, computing servers, WIFI hotspots and routers, or even network attached storage devices. And the support of so many heterogeneous device types has likely contributed to the large scale of edge nodes. Also, the ecosystem of OECPs involve a complicated interactions among three types of participants: edge node operators, OECP providers, and tenants of edge computing tasks. Among these participants, edge node operators transform their devices into edge nodes and contribute computing/bandwidth resources for revenue. Then, ECP providers are responsible for coordinating edge nodes and distributing edge computing tasks across edge nodes, while edge computing tenants subscribe to OECPs so as to deploy edge computing tasks.
@@
Another security risk is the exposure of long-term and cross-edge credentials to potential attackers. As edge nodes hosting content delivery tasks can serve as TLS servers, we find that edge nodes across platforms tend to share and locally store the same set of long-term TLS credentials (TLS private keys), which renders a non-negligible surface for MITM attacks against TLS traffic. Particularly, as long as one of these edge nodes was by-default malicious or compromised, TLS credentials for all these edge nodes would be leaked to the attackers, and transparent MITM attacks can be conducted. On the other hand, when serving as TLS clients (e.g., pulling control instructions from an edge server), edge nodes across the two OECPs fail to validate the server-side TLS certificates for part if not all of the TLS traffic flows. The negative impact of such validation failures can be well magnified by the large scale of edge nodes as well as the critical roles of these traffic flows in control-plane edge-to-server communication. Other security/privacy issues we have located include the insufficient confinement for co-located edge computing tasks, the extensive leakage of end-users’ information (e.g., the device type, IP addresses and plaintext content), and the recruitment of edge nodes without verifiable authorization from device owners, etc. We have responsibly disclosed these security findings to relevant parties including OECP operators and CDN services. By this writing, OneThingCloud and Xingyu CDN have fully acknowledged our reports and are working to fix these issues. Besides, we have also provided several recommendations to address these security risks, e.g., the Tor-like edge-based routing protocol to mitigate the privacy risks against end users. For more details, please refer to §VI.
Ethical considerations. We take ethics seriously and have carefully designed our methodology to avoid any ethical issues. Firstly, while collecting and measuring edge nodes and edge traffic, we conduct only statistical analysis to reveal scale and activity categories, avoiding any attempts to look into the content of plaintext traffic. Then, when evaluating identified security risks, we carefully design the experiments to avoid disrupting any content delivery activities. Also, all the identified security risks have been responsibly reported to the related parties, along with full acknowledgement received from OneThingCloud (§VI). On the other hand, we sought an IRB review before conducting this study, but unfortunately found out that the IRB at our institution has a limited scope of reviewing biological/medical studies and has yet to be capable of reviewing applications from other domains, including cybersecurity. Instead, we sought ethics reviews from multiple external cybersecurity researchers and have amended our methodology design by following their feedbacks.
@@ II Preliminaries
III Collecting and Analyzing Edge Activities
@@
III-A The Edge Node Deployment Framework
Despite recruiting third-party edge nodes, the two OECPs under our study are indeed blackboxes in many aspects. Particularly, little information is publicly available about their technical mechanisms, landscape of edge nodes, as well as their real-world usage. And One way to harvest this knowledge is to join both platforms as edge computing nodes. Thus, a framework is designed and implemented to achieve this goal. In a nutshell, this framework allows us to automatically deploy nodes across OECPs, capture their activities (e.g., network traffic), and backup the resulting measurement data.
When designing this framework, the first decision we should make is about what node deployment option to choose, this is because both TipTime and OneThingCloud have offered multiple options for deploying their edge nodes, so as to accommodate for as many categories of devices as possible. Specifically, Both of them provide edge node payload as firmware (mainly for router devices), docker images, Android apps while OneThingCloud even further supports running an edge node as a plugin of a network attached device. In our deployment framework, we choose the option of docker images since it allows quick and scalable deployment across heterogeneous devices. One thing to note, given observations and findings distilled from docker-based edge node deployment, it is intuitive to wonder whether they are applicable to edge nodes deployed through other options (e.g., firmware), which we have confirmed through small-scaled and short-period edge node deployment using options other than docker images.
Then, to trigger effective edge tasks, multiple resource requirements should be further satisfied. First, varied across OECPs, content delivery is carried out either through TCP or UDP flows. To support UDP-based content delivery, the edge node should at least be configured with an open NAT so that it can serve incoming UDP packets, e.g., via STUN-based NAT traversal. Furthermore, to support TCP-based content delivery, the edge node should be able to listen to a public IP address and accept incoming TCP connections. To satisfy both, we configure each edge container with the host network type, which allows it to have full access to the network of the host, a virtual machine deployed in a public cloud with static public IP addresses attached. Also, as illustrated in official tutorials [22, 23], there is a minimum disk storage requirement which varies across OECPs. Specifically, OneThingCloud requires a minimum disk storage of 50GB in 2022 and 200GB in 2023, while it is 32GB for edge nodes of TipTime. Thus, when first deploying edge nodes in 2022, we configured each edge node with 100GB disk storage, 1 virtual CPT cores, 1GB memory, and 20Mbps downstream/upstream bandwidth. Then, as OneThingCloud upgraded the minimum disk storage requirement from 50GB to 200GB in 2023, we updated the configuration as well. Besides, a runtime monitoring module is further equipped along the docker container of each edge node so as to capture its activities and resource usage. Particularly, given a running edge node, its resource consumption will be snapshotted every minute through the docker stats command. Also, tcpdump is instructed to capture its raw network traffic. All these runtime logs will be transmitted to a central backup server on a daily base and get cleaned up locally to free up disk space.
@@ III-A The Edge Node Deployment Framework
In summary, our deployment has involved 378 node-days (equivalent to running a single edge node for 378 days), which has generated 57,547,905 traffic flows and 6990.94 GB network traffic. Also, during the deployment, our edge nodes have communicated with 1,210,761 distinct IP addresses, which are widely distributed in 1,016 autonomous systems and 105 countries. More details of this traffic dataset will be presented in §IV along with an in-depth analysis of its underlying activities which turn out to be content delivery activities involving a diverse set of CDN services and content providers.
III-B The Edge Traffic Analyzer
@@
Given the large volume of edge traffic, we would like to figure out what purpose each traffic flow is intended for, reveal what remote parties have communicated with our edge nodes, and ultimately understand what edge computing activities have been conducted and through what kinds of network protocols. We pursue these tasks through a combination of manual analysis and automatic measurements.In a nutshell, the manual analysis allows us to gain qualitative knowledge such as the categories of edge traffic flows, and the signatures to associate traffic flows with different categories or distinct remote parties. On the other hand, automatic measurements are designed to generate quantitative measurement results, e.g., the volume and shares of different traffic categories, and the distribution of traffic flows over different remote parties, etc. Figure 1 illustrates how the edge traffic analyzer works, for which, the details are provided below.
Manual analysis. During our manual analysis, Wireshark was utilized to view the raw network traffic. However, it turns out to be non-trivial to locate the purposes of each traffic flow or the real-world remote parties underpinning them, especially when a traffic flow is encrypted. Fortunately, we have identified several effective tricks to conquer this challenge. One is to utilize the correlation between traffic flows especially the correlation between plaintext traffic flows and encrypted ones. For instance, as revealed by our manual study, in edge nodes of TipTime, the CDN module from YunFan will be deployed as the by-default edge task. And the CDN module periodically queries the control server regarding what content to cache and where to pull the respective content payload. Once the respective content payload(e.g., an Android APK file) is cached, a log message will be sent to the logging server of the same CDN. These traffic flows are causally related. Then, since the last traffic flow for logging is in plaintext, it can thus help us to infer the semantics of the first two traffic flows which are encrypted.
Leveraging a combination of these tricks and techniques, we have gained with high confidence a good understanding of both OECPs with regards to their technical mechanisms (i.e., network protocols), edge computing tasks, and the signatures to group traffic flows. Particularly, all edge computing tasks observed in our edge nodes turn out to be content delivery tasks that involve various CDN services and content providers. Also, a typical edge node tend to concurrently host multiple content delivery tasks that belong to different parties. More details will be presented in §IV.
@@
Then, when it comes to assigning a traffic flow to different categories, aforementioned manual analysis has identified signatures to group traffic into the following 4 main categories: edge-wide control-plane, CDN-specific control-plane, content caching, and content delivery. Among these categories, edge-wide control-plane traffic encompasses the traffic flows between edge nodes and control servers of the respective OECP, which are aimed to fulfill goals including pulling control instructions, pushing logging data, and downloading executable payloads of edge computing tasks, etc. Then, as all edge computing tasks that we have observed are content delivery tasks, the second traffic category is CDN-specific control-plane traffic between an on-edge CDN module and the respective remote CDN control server. Then, across content delivery tasks, two more categories of traffic flows are observed to fulfill content delivery. One is the flows to cache content from either upstream content servers or counterpart CDN nodes, which we name as the content caching traffic. The other category involves traffic flows that are intended to delivery content payloads to either end devices (e.g., viewers of videos) or counterpart CDN nodes, and we thus call it the content delivery traffic. Then, since multiple CDN modules can be co-deployed on the same edge node, traffic flows of the same CDN-relevant category can be further correlated with a specific CDN service (e.g., YunFan), or even a specific content provider (e.g., Douyin). In a nutshell, the traffic signatures identified through manual analysis enable us to group all edge traffic flows into these four main categories as well as correlating most content delivery/caching traffic flows with respective CDN service and content provider. For instance, we consider a content delivery traffic flow belongs to YunFan CDN and is used to delivery content to end users of the Douyin platform as long as its FQDN matches the pattern of *yf*.free-lbv6.idouyinvod.com.
IV The Ecosystem
@@
Leveraging aforementioned methodology, we have gained a deep understanding for the ecosystem of open edge computing. As shown in Figure 2, multiple parties have participated in this ecosystem. First of all, open edge computing platforms (OECPs) build up the pool of edge nodes through recruiting third-party devices, e.g, personal computers, mobile phones, WIFI hotspots. As the return, these device owners, i.e., edge node operators, can earn a revenue that is typically proportional to the volume of resource consumption, e.g., bandwidth, computing, and storage resources. Then, Given the computing/bandwidth resources available in these edge nodes, various content delivery services (CDNs) are the main consumers which deploy their content delivery modules on edge nodes and transform these edge nodes into content delivery servers (i.e., CDN nodes). One more indirect participant are the content providers (CPs), e.g., video streaming platforms, or cloud storage services. Instead of directly collaborating with OECPs, these CPs subscribe to one or more CDN services, which in turn, instruct edge nodes via the on-edge CDN modules to pull content from content providers, cache content locally, and deliver the content to end users on demand. Below, we provide more detailed characterization for these participants with regards to their scale, distribution, evolution, etc. And the resulting understanding of this ecosystem serves as an important prior before we can well profile the impact of the inherent security risks (§V).
IV-A Open Edge Computing Platforms
@@
Technical mechanisms. In a nutshell, a typical edge node is responsible to orchestrate and monitor edge computing tasks. Once deployed and started, the edge node communicates with the OECP control server, which in turn, returns a list of edge computing tasks along with the instructions for deploying these tasks. Then, to deploy an edge computing task, the edge node downloads the necessary executable payloads and jump-starts the respective edge computing task. When executing edge computing tasks, the edge node periodically sends logs (e.g., resource consumption statistics) to the control server for diagnosis and billing. Such a paradigm is applicable to edge nodes of both OECPs. One thing to note, edge computing tasks deployed on the same edge node can belong to different real-world parties, most of which are independent from and collaborate with the OECPs. However, the two OECPs differ in how they isolate these co-located edge computing tasks. Briefly, a OneThingCloud edge node runs edge computing tasks as separate containers, whereas a TipTime edge node runs tasks under the root user and enforces no confinement for them, which obviously incurs non-negligible security concerns, as further elaborated in §V.
Edge computing tasks. Regarding the types of edge computing tasks, all tasks observed in our deployment are CDN tasks. One explanation is that peer-to-peer CDN (PCDN) is in great demand on the China CDN market. However, variance still exists in their task operator and content types. For instance, TipTime edge nodes under our deployment have been used to execute content delivery tasks from two CDN services, namely, YunFan CDN and Wangsu CDN. Also, the CDN modules of YunFan have been observed to deliver content payloads for 7 different and highly popular content providers, e.g., Kuishou and Douyin in the domain of short video streaming. More details about CDN services and content providers will be elaborated in §IV-C.
IV-B Edge Computing Nodes

### v3

@@ 1 Introduction
However, these OECPs are black boxes since little is publicly available regarding their technical mechanisms or real-world usage. Furthermore, it is intuitive to wonder what practical security/privacy implications these OECPs may render, especially considering the claimed extensive adoption by online services. For instance, our preliminary investigation reveals that none of these OECPs enforce strict vetting for edge node operators, and any third party can turn their devices into edge nodes, which can introduce byzantine or even malicious edge nodes. It also breaks the traditional threat model for edge computing wherein edge nodes are fully controlled by the ECP and are usually considered as trusted. To clear away the dense mist surrounding OECPs, we conduct an extensive security study on two representative ones, namely, TipTime [10] and OneThingCloud [9]. As elaborated in §2, these two stand out in a comparative analysis of all identified OECPs. Our study would not be possible without the design and implementation of a set of automatic or semi-automatic tools. Particularly, to capture edge activities and edge traffic, a deployment framework is designed to deploy edge nodes across OECPs at different locations and capture their edge traffic. Then, given edge traffic of multiple TBs, an efficient edge traffic analyzer is built up to concurrently process edge traffic flows, group them into pre-defined traffic categories, as well as correlating them with various real-world parties. Also, a security testbed is carefully designed to empirically evaluate security risks of OECPs while avoiding interfering with real-world edge computing activities. Applying these tools leads to new understandings of OECPs with regards to their technical mechanisms, the landscape of edge nodes, real-world usage, as well as security and privacy risks.
@@
Our key findings are highlighted as below. First of all, Both OECPs share many properties in their technical mechanism and service model. Briefly, both enable the co-deployment of different edge computing tasks on a single edge node and each edge computing task runs like an independent kingdom inside the edge node. Also, when it comes to edge node recruitment, both support a variety of computing devices including mobile phones, desktop/laptop computers, computing servers, WIFI hotspots and routers, or even network attached storage devices. And the support of so many heterogeneous device types has likely contributed to the large scale of edge nodes. Also, the ecosystem of OECPs involve a complicated interactions among three types of participants: edge node operators, OECP providers, and tenants of edge computing tasks. Among these participants, edge node operators transform their devices into edge nodes and contribute computing/bandwidth resources for revenue. Then, ECP providers are responsible for coordinating edge nodes and distributing edge computing tasks across edge nodes, while edge computing tenants subscribe to OECPs so as to deploy edge computing tasks.
@@
Another security risk is the exposure of long-term and cross-edge credentials to potential attackers. As edge nodes hosting content delivery tasks can serve as TLS servers, we find that edge nodes across platforms tend to share and locally store the same set of long-term TLS credentials (TLS private keys), which renders a non-negligible surface for MITM attacks against TLS traffic. Particularly, as long as one of these edge nodes was by-default malicious or compromised, TLS credentials for all these edge nodes would be leaked to the attackers, and transparent MITM attacks can be conducted. On the other hand, when serving as TLS clients (e.g., pulling control instructions from an edge server), edge nodes across the two OECPs fail to validate the server-side TLS certificates for part if not all of the TLS traffic flows. The negative impact of such validation failures can be well magnified by the large scale of edge nodes as well as the critical roles of these traffic flows in control-plane edge-to-server communication. Other security/privacy issues we have located include the insufficient confinement for co-located edge computing tasks, the extensive leakage of end-users’ information (e.g., the device type, IP addresses and plaintext content), and the recruitment of edge nodes without verifiable authorization from device owners, etc. We have responsibly disclosed these security findings to relevant parties including OECP operators and CDN services. By this writing, OneThingCloud and Xingyu CDN have fully acknowledged our reports and are working to fix these issues. Besides, we have also provided several recommendations to address these security risks, e.g., the Tor-like edge-based routing protocol to mitigate the privacy risks against end users. For more details, please refer to §5.7.
Ethical considerations. We take ethics seriously and have carefully designed our methodology to avoid any ethical issues. Firstly, while collecting and measuring edge nodes and edge traffic, we conduct only statistical analysis to reveal scale and activity categories, avoiding any attempts to look into the content of plaintext traffic. Then, when evaluating identified security risks, we carefully design the experiments to avoid disrupting any content delivery activities. Also, all the identified security risks have been responsibly reported to the related parties, along with full acknowledgement received from OneThingCloud (§5.7). On the other hand, we sought an IRB review before conducting this study, but unfortunately found out that the IRB at our institution has a limited scope of reviewing biological/medical studies and has yet to be capable of reviewing applications from other domains, including cybersecurity. Instead, we sought ethics reviews from multiple external cybersecurity researchers and have amended our methodology design by following their feedbacks.
@@ 2 Preliminaries
3 Collecting and Analyzing Edge Activities
@@
3.1 The Edge Node Deployment Framework
In a nutshell, this framework allows us to automatically deploy nodes across OECPs, capture their activities (e.g., network traffic), and backup the resulting measurement data.
Both TipTime and OneThingCloud have offered multiple options for deploying their edge nodes, and the full list of deployment optionscan be found in Appendix A. In our deployment framework, we choose the option of Docker images since it allows quick and scalable deployment across heterogeneous devices. To support incoming connections that are either TCP or UDP, we configure each edge container with the host network type, which allows it to have full access to the network of the host, a virtual machine deployed in a public cloud with static public IP addresses attached. Also, as illustrated in official tutorials [22, 23], there is a minimum disk storage requirement which varies across OECPs. Specifically, OneThingCloud requires a minimum disk storage of 50GB in 2022 and 200GB in 2023, while it is 32GB for edge nodes of TipTime. Thus, when first deploying edge nodes in 2022, we configured each edge node with 100GB disk storage, 1 virtual CPT cores, 1GB memory, and 20Mbps downstream/upstream bandwidth. Then, as OneThingCloud upgraded the minimum disk storage requirement from 50GB to 200GB in 2023, we updated the configuration as well. Also, tcpdump is instructed to capture its raw network traffic.
@@ 3.1 The Edge Node Deployment Framework
In summary, our deployment has involved 378 node-days (equivalent to running a single edge node for 378 days), which has generated 57,547,905 traffic flows and 6990.94 GB network traffic. Also, during the deployment, our edge nodes have communicated with 1,210,761 distinct IP addresses, which are widely distributed in 1,016 autonomous systems and 105 countries. More details of this traffic dataset will be presented in §4 along with an in-depth analysis of its underlying activities which turn out to be content delivery activities involving a diverse set of CDN services and content providers.
3.2 The Edge Traffic Analyzer
@@
Given the large volume of edge traffic, we would like to figure out what purpose each traffic flow is intended for, reveal what remote parties have communicated with our edge nodes, and ultimately understand what edge computing activities have been conducted and through what kinds of network protocols. We pursue these tasks through a combination of manual analysis and automatic measurements. As shown in Figure 1, the manual analysis allows us to gain qualitative knowledge such as the categories of edge traffic flows, and the signatures to associate traffic flows with different categories or distinct remote parties. On the other hand, automatic measurements are designed to generate quantitative measurement results, e.g., the volume and shares of different traffic categories.
Manual analysis. During our manual analysis, Wireshark was utilized to view the raw network traffic. However, it turns out to be non-trivial to locate the purposes of each traffic flow or the real-world remote parties underpinning them, especially when a traffic flow is encrypted. Fortunately, we have identified several effective tricks to conquer this challenge, including utilizing the correlation among traffic flows especially the correlation between plaintext traffic flows and encrypted ones, mapping domain names to their real-world entities by querying credible domain WHOIS dataset, and observating the difference between downstream and upstream traffic. More details can be found in Appendix A.
Leveraging a combination of these tricks and techniques, we have gained with high confidence a good understanding of both OECPs with regards to their technical mechanisms (i.e., network protocols), edge computing tasks, and the signatures to group traffic flows. Particularly, all edge computing tasks observed in our edge nodes turn out to be content delivery tasks that involve various CDN services and content providers. Also, a typical edge node tend to concurrently host multiple content delivery tasks that belong to different parties. More details will be presented in §4.
@@
Then, when it comes to assigning a traffic flow to different categories, aforementioned manual analysis has identified signatures to group traffic into the following 4 main categories: edge-wide control-plane, CDN-specific control-plane, content caching, and content delivery. Among these categories, edge-wide control-plane traffic encompasses the traffic flows between edge nodes and control servers of the respective OECP, which are aimed to fulfill goals including pulling control instructions, pushing logging data, and downloading executable payloads of edge computing tasks, etc. Then, as all edge computing tasks that we have observed are content delivery tasks, the second traffic category is CDN-specific control-plane traffic between an on-edge CDN module and the respective remote CDN control server. Then, across content delivery tasks, two more categories of traffic flows are observed to fulfill content delivery. One is the flows to cache content from either upstream content servers or counterpart CDN nodes, which we name as the content caching traffic. The other category involves traffic flows that are intended to delivery content payloads to either end devices (e.g., viewers of videos) or counterpart CDN nodes, and we thus call it the content delivery traffic. Then, since multiple CDN modules can be co-deployed on the same edge node, traffic flows of the same CDN-relevant category can be further correlated with a specific CDN service (e.g., YunFan), or even a specific content provider (e.g., Douyin). In a nutshell, the traffic signatures identified through manual analysis enable us to group each traffic flow into these four main categories as well as correlating most content delivery/caching traffic flows with respective CDN service and content provider. For instance, we consider a content delivery traffic flow belongs to YunFan CDN and is used to delivery content to end users of the Douyin platform as long as its FQDN matches the pattern of *yf*.free-lbv6.idouyinvod.com.
4 The Ecosystem
@@
Leveraging aforementioned methodology, we have gained a deep understanding for the ecosystem of open edge computing. As shown in Figure 2, multiple parties have participated in this ecosystem. First of all, open edge computing platforms (OECPs) build up the pool of edge nodes through recruiting third-party devices, e.g, personal computers, mobile phones, WIFI hotspots. As the return, these device owners, i.e., edge node operators, can earn a revenue that is typically proportional to the volume of resource consumption, e.g., bandwidth, computing, and storage resources. Then, Given the computing/bandwidth resources available in these edge nodes, various content delivery services (CDNs) are the main consumers which deploy their content delivery modules on edge nodes and transform these edge nodes into content delivery servers (i.e., CDN nodes). One more indirect participant are the content providers (CPs), e.g., video streaming platforms, or cloud storage services. Instead of directly collaborating with OECPs, these CPs subscribe to one or more CDN services, which in turn, instruct edge nodes via the on-edge CDN modules to pull content from content providers, cache content locally, and deliver the content to end users on demand. Below, we provide more detailed characterization for these participants with regards to their scale, distribution, evolution, etc. And the resulting understanding of this ecosystem serves as an important prior before we can well profile the impact of the inherent security risks (§5).
4.1 Open Edge Computing Platforms
@@
Technical mechanisms. In a nutshell, a typical edge node is responsible to orchestrate and monitor edge computing tasks. Once deployed and started, the edge node communicates with the OECP control server, which in turn, returns a list of edge computing tasks along with the instructions for deploying these tasks. Then, to deploy an edge computing task, the edge node downloads the necessary executable payloads and jump-starts the respective edge computing task. When executing edge computing tasks, the edge node periodically sends logs (e.g., resource consumption statistics) to the control server for diagnosis and billing. Such a paradigm is applicable to edge nodes of both OECPs. One thing to note, edge computing tasks deployed on the same edge node can belong to different real-world parties, most of which are independent from and collaborate with the OECPs. However, the two OECPs differ in how they isolate these co-located edge computing tasks. Briefly, a OneThingCloud edge node runs edge computing tasks as separate containers, whereas a TipTime edge node runs tasks under the root user and enforces no confinement for them, which obviously incurs non-negligible security concerns, as further elaborated in §5.
Edge computing tasks. Regarding the types of edge computing tasks, all tasks observed in our deployment are CDN tasks. One explanation is that peer-to-peer CDN (PCDN) is in great demand on the China CDN market. However, variance still exists in their task operator and content types. For instance, TipTime edge nodes under our deployment have been used to execute content delivery tasks from two CDN services, namely, YunFan CDN and Wangsu CDN. Also, the CDN modules of YunFan have been observed to deliver content payloads for 7 different and highly popular content providers, e.g., Kuishou and Douyin in the domain of short video streaming. More details about CDN services and content providers will be elaborated in §4.3.
4.2 Edge Computing Nodes